OpenStudy (princesssleelee):

Which of the following is a polynomial with roots negative square root of 3, square root of 3, and −2? x3 − 2x2 − 3x + 6 x3 + 2x2 − 3x − 6 x3 − 3x2 − 5x + 15 x3 + 3x2 − 5x − 15

OpenStudy (amorfide):

You are given the three roots \[x=\sqrt{3}, x=-\sqrt{3}, x=-2\] So you can just substitute these into the given polynomials, if the answer is 0 then it is a root of that polynomial for example \[x^3 - 2x^2 -3x +6\] let \[x=\sqrt{3}\] then \[\sqrt{3}^{3} -2(\sqrt{3})^2 -3\sqrt{3} + 6 = 3\sqrt{3} - 6 - 3\sqrt{3} + 6 = 0\] then repeat by letting x equal the other two roots, and continue to substitute them into all of the given polynomials, until a polynomial gives 0 as the answer for all three roots

OpenStudy (princesssleelee):

I dont quite understand..

OpenStudy (princesssleelee):

@amorfide

OpenStudy (mathmale):

I believe y our best bet is to form the three factors of this polynomial and to multiply them together. That will most positively help you choose the correct answer. If x=-2 is a root, then (x+2) is a factor. See that sign change? If \[x=\sqrt{3}\]

OpenStudy (mathmale):

is a root, then \[(x-\sqrt{3}) \]

OpenStudy (princesssleelee):

I am so lost.

OpenStudy (mathmale):

is a factor. If negative sqrt 3 is a root, \[x=-\sqrt{3}\]

OpenStudy (mathmale):

is a factor. If neg sqrt (3) is a root, what is the 3rd factor?

OpenStudy (mathmale):

I'm sorry you find this to be difficult to grasp. Could you possibly ask questions? Let me work out one sample problem for you. Given that x=2 and x=3 are roots of a quadratic equation, write the actual quadratic.

OpenStudy (princesssleelee):

Could you work this one out lol, theres another just like it that i have to do using the same numbers

OpenStudy (mathmale):

Take that x=2 and write the factor (x-2); take that x=3 and write the factor (x-3). You really do have to know and undrstand this fact.

OpenStudy (mathmale):

What is the polynomial? Use the FOIL method to multiply: (x-2)(x-3). Can you do that?

OpenStudy (mathmale):

First terms: x times x = x^2 ("x squared")

OpenStudy (mathmale):

Outer terms: (x)(3) = 3x

OpenStudy (mathmale):

Inner terms: (-2x)(-3x) = ?

OpenStudy (princesssleelee):

x^2-3x-2x+6

OpenStudy (mathmale):

Yes, and then combine the 2 middle terms. What do you get?

OpenStudy (princesssleelee):

x2-5x+6

OpenStudy (mathmale):

Perfect, except that you should write the first term as x^2, or as \[x^2.\]

OpenStudy (princesssleelee):

correct

OpenStudy (mathmale):

What is your final answer? In other words, where is the 3rd order polynomial you've created?

OpenStudy (mathmale):

Looks like I may have misled you. If you have roots POS Sqrt(3), NEG Sqrt(3), and -2, write the appropriate factors:\[(x+\sqrt{2})(x-\sqrt{2})(x+2)\]

OpenStudy (mathmale):

Sorry for the extra work. Pls multiply out these 3 factors and simplify.\[(x-\sqrt{3})(x+\sqrt{3})(x+2).\]

OpenStudy (mathmale):

Important hint:

OpenStudy (mathmale):

(a-b)(a+b)=a^2-b^2